OpenStudy (anonymous):

How many solutions can be found for the equation 4x = 4x? Zero One Two Infinitely many

OpenStudy (michele_laino):

hint: the left side is identical to the right side for every value of x

OpenStudy (anonymous):

so zero

OpenStudy (michele_laino):

if we replace x with 1 we have: 4*1=4*1 4=4

OpenStudy (michele_laino):

if we replace x with 2, we have: 4*2=4*2 8=8

OpenStudy (michele_laino):

and so on...

OpenStudy (anonymous):

oh ok its D

OpenStudy (michele_laino):

that's right!